# Supersonic Flow Past an Obstacle in a Quasi-Two-Dimensional Lee–Huang–Yang Quantum Fluid

### v1

@@ I Introduction
where $U_{ext}$ represents the external potential, which can account for effects such as traps and obstacles, among others. The parameters $g_{LHY}$ and $m$ denote the three-dimensional inter-particle interaction constant and the atomic mass, respectively. The quantum fluid governed by the LHY equation exhibits both nonlinear and dispersive effects. These properties enable the propagation of various atomic excitations, including solitons, vortices, dispersive shock waves, and linear radiation. As in the case of the GPE, a supersonic flow past an obstacle in the LHY fluid is expected to generate stationary coherent wave patterns satisfying the Mach-Cherenkov-Landau resonant radiation condition. These wakes are referred to as Kelvin-like wakes due to their similarity to Lord Kelvin’s water wave structure generated by a moving ship (commonly known as the Kelvin wake pattern) [7, 8, 9, 10]. More specifically, for supersonic flow velocities, the GPE predicts two distinct wave structures: oblique dark solitons inside the Mach cone and linear radiation outside it. In the case of the GPE, these structures have been extensively studied, experimentally [11, 12] and through theoretical investigations, particularly regarding linear radiation [13, 14] and oblique dark soliton wakes [15, 16]. However, for Eq. (1), this problem remains unsolved.
@@ II Dimensional Reduction and the Quasi-2D Hydrodynamical Model
We will consider the model of quasi-2D BEC tightly confined in the $z$-direction by the external potential, so that the condensate cloud acquires a “pancake” configuration. This means the harmonic potential corresponds to much greater frequency $\omega_{z}$ in $z$ of atomic motion direction than the frequency $\omega_{\perp}$ of oscillations in the “pancake” plane, $\omega_{z}/\omega_{\perp}\gg 1$. Consequently, the atoms condense into the ground state $\phi(z)$ of the motions in the $z$-direction for which the weak interatomic interaction can be neglected, whereas the atomic motion in the $(x,y)$-plane is governed by the Gross-Pitaevskii equation with account of the interatomic interaction. If number of atoms $N$ in the condensate is large enough, so that the radius $R_{0}$ of the pancake-shape density distribution is much greater than the characteristic harmonic length $a_{z}$ in the axial direction, then we can factorize the 3D condensate wave function in the following way,
@@
$\phi(z)=\frac{1}{\pi^{1/4}a_{z}^{1/4}}e^{-\left(\frac{z^{2}}{2a_{z}}\right)}$ (3)
@@
where $\xi_{0}=\hbar/\sqrt{m|\Psi_{0}|^{3}g^{\prime}_{LHY}}$, $\tau=\hbar/|\Psi_{0}|^{3}g^{\prime}_{LHY}$ and $\Psi_{0}=1/|\Psi_{0}|^{3}g^{\prime}_{LHY}$, we transform Eq. (4) to non-dimensional variables. As a result, we arrive at the equation
@@
[…]
@@ III Linear radiation in LHY fluid flowing past an obstacle
$\begin{split}&\frac{\partial\rho^{\prime}}{\partial t}+u_{0}\frac{\partial\rho^{\prime}}{\partial x}+\rho_{0}\left(\frac{\partial u^{\prime}}{\partial x}+\frac{\partial v^{\prime}}{\partial y}\right)=0,\\ &\frac{\partial u^{\prime}}{\partial t}+u_{0}\frac{\partial u^{\prime}}{\partial x}+\frac{3}{2}\sqrt{\rho_{0}}\frac{\partial\rho^{\prime}}{\partial x}=\\ &\hskip 56.9055pt\frac{1}{4\rho_{0}}\left(\frac{\partial^{3}\rho^{\prime}}{\partial x^{3}}+\frac{\partial^{3}\rho^{\prime}}{\partial x\partial y^{2}}\right),\\ &\frac{\partial v^{\prime}}{\partial t}+u_{0}\frac{\partial v^{\prime}}{\partial x}+\frac{3}{2}\sqrt{\rho_{0}}\frac{\partial\rho^{\prime}}{\partial y}=\\ &\hskip 56.9055pt\frac{1}{4\rho_{0}}\left(\frac{\partial^{3}\rho^{\prime}}{\partial y^{3}}+\frac{\partial^{3}\rho^{\prime}}{\partial y\partial x^{2}}\right).\\ \end{split}$ (11)
@@
is called the ‘sound velocity’. Introduction of this characteristic flow velocity allows one to rewrite the dispersion relation, making the transformation in dimensionless variables
@@
where $M$ denotes the Mach number and $k$ is the modulus of the wave vector $\mathbf{k}=(k_{x},k_{y})$. This form of the dispersion relation coincides with that for linear waves in usual BEC, whose dynamics is described by the standard GPE (see, e.g. Ref. [13]), so we can use the method of Refs. [8, 13] for finding the Kelvin-like wake pattern.
@@
$\mathbf{k}=\nabla\theta,\quad G(k_{y},k_{x})=\omega=0.$ (17)
@@
At last, according to the definition of the angles in Fig. 1, we have $y/x=\tan{\xi}$ ), and then straightforward algebra (see Ref. [13]) yields
$\tan{\xi}=\frac{\left(2M^{2}-1-\tan^{2}\eta\right)\tan\eta}{\left(M^{2}+1\right)\tan^{2}\eta-\left(M^{2}-1\right)}.$ (23)
@@
[…]
Since the wave vector is a gradient of the phase, the path of the integration can be chosen arbitrarily, and it is convenient integrate along the line with $\xi=\mathrm{const}$, so we get
@@
$-\arccos\left(\frac{1}{M}\right)\leq\eta\leq\arccos\left(\frac{1}{M}\right).$ (31)
@@ IV.1 DARK SOLITONS IN A 1D LHY FLUID
The soliton solution can be found following the general theory developed in Ref. [18, 15]. The 1D LHY equation has the form
@@
[…]
@@ IV.2 OBLIQUE DARK SOLITON IN A QUASI-2D SYSTEM
$x^{\prime}=x\cos\alpha-y\sin\alpha\text{, \quad and \quad}y^{\prime}=x\sin\alpha-y\cos\alpha.$ (43)
@@ V Numerical experiment and comparison with analytical results
Figure 3: Comparison of the wave crest position calculated analytically (see (III)) and the numerical simulations. The analytical solution is represented by the dashed line and the results correspond to $M=2$. The simulation was conducted in a quasi-2D system for an impenetrable obstacle with a small radius $R=1$ and evolution time $t=20$.
@@ V.2 Numerical Method
$\begin{split}i\partialderivative{\varphi}{t^{\prime}}=&-\frac{1}{2}\quantity(\frac{\partial^{2}\varphi}{\partial x^{\prime 2}}+\frac{\partial^{2}\varphi}{\partial y^{\prime 2}})+iv\partialderivative{\varphi}{x^{\prime}}-\rho_{0}\varphi\\ &+U_{obs}(x^{\prime},y^{\prime})\varphi+\absolutevalue{\varphi}^{3}\varphi,\end{split}$ (46)
@@
Given the two-dimensional nature of the system, we use the Peaceman-Rachford (PR) method [20] for spatial derivatives, combined with the split-step operator method for the ordinary terms. The PR method is well known for being unconditionally stable when applied exclusively to the dispersion terms. In our implementation, it proved to be effective for evolving the system in time while maintaining numerical accuracy and stability. Additionally, suitable discretizations techniques were applied to ensure that both the boundary conditions and the presence of the obstacle were accurately incorporated into the numerical framework. The typical spatial and temporal discretizations employed are $\Delta x=\Delta y=0.20$ and $\Delta t=0.001$, respectively.
V.3 LINEAR RADIATION
As is shown in Fig. 3, wave crests of linear radiation can be calculated analytically with good precision from Eq. (III). Moreover, Eq. (III) predicts that just ahead of the obstacle, in the limit $y\rightarrow 0$, the wavelength vanishes as the flow velocity $|u_{0}|$ increases. This becomes clear by assuming $\psi\rightarrow 0$ and expanding Eq. (III) in a Taylor series. This calculation results in the system:
$\begin{split}&x\approx\frac{\theta}{2\sqrt{u_{0}^{2}-c_{0}^{2}}}+\frac{\theta(c_{0}^{2}-5u_{0}^{2})}{\left(u_{0}^{2}-c_{0}^{2}\right)^{3/2}}\psi^{2},\\ &y\approx\frac{\theta\left(2u_{0}^{2}-c_{0}^{2}\right)}{2\left(u_{0}^{2}-c_{0}^{2}\right)^{3/2}}\psi,\end{split}$ (47)
@@ V.3 LINEAR RADIATION
which decreases as a function of the velocity $u_{0}$. This “shortening” of the wavelength can be seen when we compare the wave patterns for $M=2$ in Fig. 3 with the other one for $M=5$ in Fig. 4. If we increase $M$ even more, then we observe in numerical simulations a “continuous” density distribution in the region of for small $\psi$. This means that the wavelength becomes here smaller than the size of the obstacle, so the interference of generated harmonics becomes destructive and it does not lead to formation of a regular wake wave pattern.
@@
Figure 4: Simulation performed for an uniform quasi-2D system with an impenetrable obstacle of radius $R=1$. The numerical experiment was conducted for $M=5$ and $t=30$, where the emergence of a dark soliton can be observed.
V.4 OBLIQUE DARK SOLITON
@@
The oblique dark solitons produced by the LHY fluid inside the Mach cone attains a stationary form only when the critical flow velocity is reached. Below this critical velocity, the soliton is unstable and, similar to the case of the Gross-Pitaevskii equation, it breaks into vortex-antivortex pairs due to the snake instability. A discussion of this instability in the context of the Gross-Pitaevskii equation can be found in Ref. [21] and qualitatively similar behavior is expected for the case of LHY dynamics. An example of a non-stationary street of vortices instead of a stationary oblique soliton is shown in Fig. 3. Above the critical velocity $M>M_{c}$, the oblique dark soliton is formed, as is shown in Fig. 4. A rough estimate based on our numerical results suggests that the critical Mach number for soliton formation lies in the range $3\lesssim M_{c}\lesssim 3.5$. The agreement between the analytical solution for the soliton’s density profile given by Eq. (45) and the numerical profile extracted from Fig. 4 at $x=-80$ is very good, as is demonstrated in Fig. 5.
@@
Figure 5: Analytical and numerical profiles of the dark soliton for $M=5$ and $x=-80$. The analytical results were obtained using Eq. (45), where the value of the angle $\alpha=85.035^{\circ}$ was found from the numerical calculations.
@@ VI Conclusion
From an experimental perspective, the supersonic flow past an impenetrable barrier, as modeled in this work, offers a useful approximation for BEC systems past strong laser beams. The features identified here—such as oblique solitons and linear wakes—are potentially observable in current experimental setups, particularly in quantum droplets or strongly interacting mixtures where LHY corrections play a dominant role. As a direction for future research, we propose extending this framework to penetrable, weak obstacles, where the local flow velocity varies along the soliton line. This generalization is relevant for describing experiments in which a weak laser beam moves through a dilute BEC [22]. This framework may also serve as a tool to measure and analyze experimentally critical velocities associated with the emission of collective excitations as proposed in Ref. [23].

### v2

@@ I Introduction
where $U_{ext}$ represents the external potential, which can account for effects such as traps and obstacles, among others. The parameters $g_{LHY}$ and $m$ denote the three-dimensional inter-particle interaction constant and the atomic mass, respectively. The quantum fluid governed by the LHY equation exhibits both nonlinear and dispersive effects. These properties enable the propagation of various atomic excitations, including solitons, vortices, dispersive shock waves, and linear radiation. As in the case of the GPE, a supersonic flow past an obstacle in the LHY fluid is expected to generate stationary coherent wave patterns satisfying the Mach-Cherenkov-Landau resonant radiation condition. These wakes are referred to as Kelvin-like wakes due to their similarity to Lord Kelvin’s water wave structure generated by a moving ship (commonly known as the Kelvin wake pattern) [7, 8, 9, 10, 11]. More specifically, for supersonic flow velocities, the GPE predicts two distinct wave structures: oblique dark solitons inside the Mach cone and linear radiation outside it. In the case of the GPE, these structures have been extensively studied, experimentally [12, 13] and through theoretical investigations, particularly regarding linear radiation [14, 15] and oblique dark soliton wakes [16, 17]. However, for Eq. (1), this problem remains unsolved.
@@ II Dimensional Reduction and the Quasi-2D Hydrodynamical Model
In experimental setups, it is not possible to realize truly low-dimensional systems. Instead, an effective low-dimensional regime can be obtained by “squeezing” a 3D BEC, tightly confining it in one or two spatial directions using external potentials. Theoretically, beyond-mean-field energy corrections in Bose gases have been investigated within the framework of 3D–1D and 3D–2D dimensional crossovers in Refs. [18, 19, 20].
In our model, we construct a 3D–2D dimensional crossover by imposing strong confinement along the $z$-axis, such that the system’s dynamics are effectively restricted to the xxyy plane. This means that the harmonic potential corresponds to much greater frequency $\omega_{z}$ in $z$ of atomic motion direction than the frequency $\omega_{\perp}$ of oscillations in the “pancake” plane, $\omega_{z}/\omega_{\perp}\gg 1$. Consequently, the atoms condense into the ground state $\phi(z)$ of the motions in the $z$-direction for which the weak interatomic interaction can be neglected, whereas the atomic motion in the $(x,y)$-plane is governed by the Gross-Pitaevskii equation with account of the interatomic interaction. If number of atoms $N$ in the condensate is large enough, so that the radius $R_{0}$ of the pancake-shape density distribution is much greater than the characteristic harmonic length $a_{z}=\sqrt{\hbar/m\omega_{z}}$ in the axial direction, then we can factorize the 3D condensate wave function in the following way,
@@
$\phi(z)=\frac{1}{\pi^{1/4}a_{z}^{1/2}}e^{-\left(\frac{z^{2}}{2a_{z}^{2}}\right)}$ (3)
@@
where $\xi_{0}=\hbar/\sqrt{m|\Psi_{0}|^{3}g^{\prime}_{LHY}}$, and $\tau=\hbar/|\Psi_{0}|^{3}g^{\prime}_{LHY}$, we transform Eq. (4) to non-dimensional variables. As a result, we arrive at the equation
@@
$\Psi(\mathbf{r},t)=\sqrt{\rho(\mathbf{r},t)}e^{i\int\mathbf{u}(\mathbf{r})\,d\mathbf{r}}.$ (8)
@@ III Linear radiation in LHY fluid flowing past an obstacle
[…]
@@
is called the ‘sound velocity’. Introduction of this characteristic flow velocity allows one to rewrite the dispersion relation, making the transformation to dimensionless variables
@@
where $M=u_{0}/c_{0}$ denotes the Mach number and $k$ is the modulus of the wave vector $\mathbf{k}=(k_{x},k_{y})$. This form of the dispersion relation coincides with that for linear waves in usual BEC, whose dynamics is described by the standard GPE (see, e.g. Ref. [14]), so we can use the method of Refs. [8, 14] for finding the Kelvin-like wake pattern.
@@
$\mathbf{k}=\nabla\theta,\quad G(k_{x},k_{y})=\omega=0.$ (17)
@@
[…]
@@
$\theta=\int_{0}^{r}\mathbf{k}\cdot d\mathbf{r}.$ (26)
Since the wave vector is a gradient of the phase, the path of the integration can be chosen arbitrarily, and it is convenient to integrate along the line with $xi=\mathrm{const}$, so we get
@@
$\arccos\left(\frac{1}{M}\right)\leq\eta\leq\arccos\left(\frac{1}{M}\right).$ (31)
@@ IV.1 DARK SOLITONS IN A 1D LHY FLUID
A detailed discussion of the soliton solution and its stability can be found within the general theoretical framework developed in Refs. [22, 16, 23]. The 1D LHY equation has the form
@@
$\psi(\tilde{x},t)=\sqrt{\rho(\tilde{x},t)}e^{i\int\tilde{u}(\tilde{x})\,d\tilde{x}}$, (33)
@@ IV.2 OBLIQUE DARK SOLITON IN A QUASI-2D SYSTEM
[…]
@@ V.1 NUMERICAL METHOD
[…]
@@
Given the two-dimensional nature of the system, we use the Peaceman–Rachford (PR) method [25] to handle spatial derivatives, combined with the split-step operator method for the ordinary terms. The PR scheme is well known for being unconditionally stable when applied exclusively to the dispersive components of the equation. In our implementation, it proved effective for evolving the system in time while maintaining numerical accuracy and stability.
To initialize the simulation, the wavefunction at $t=0$ is set to a constant value, $\psi(x,y,0)=1$. The obstacle is modeled as an impenetrable circular barrier, incorporated through a piecewise-defined potential. Specifically, we define the potential $U(x,y)$ to be zero outside a circular region of radius $R_{0}$, and set it to a large constant $V_{0}=10^{12}$ within the region $\sqrt{x_{i}^{2}+y_{j}^{2}}\leq R_{0}$, effectively creating a hard-wall condition that enforces $\psi=0$ inside the obstacle. The typical spatial and temporal discretizations used in our simulations are $\Delta x=\Delta y=0.20$ and $\Delta t=0.001$, respectively.
Figure 3: (a) Comparison between the analytically predicted wave crest position [ Eq. (III)] and numerical simulations Eq. (46)]. The analytical result is shown as a dashed line. (b) Numerical solution for the phase. The simulation was performed in a quasi-2D system with an impenetrable obstacle of radius $R=1$, at evolution time $t=25$ and Mach number $M=2$.
V.2 LINEAR RADIATION
As is shown in Fig. 3, wave crests of linear radiation can be calculated analytically with good precision from Eq. (III). Moreover, Eq. (III) predicts that just ahead of the obstacle, in the limit $y\rightarrow 0$, the wavelength vanishes as the flow velocity $|u_{0}|$ increases. This becomes clear by assuming $\eta\rightarrow 0$ and expanding Eq. (III) in a Taylor series. This calculation results in the system:
$\begin{split}&x\approx\frac{\theta}{2\sqrt{u_{0}^{2}-c_{0}^{2}}}+\frac{\theta(c_{0}^{2}-5u_{0}^{2})}{\left(u_{0}^{2}-c_{0}^{2}\right)^{3/2}}\eta^{2},\\ &y\approx\frac{\theta\left(2u_{0}^{2}-c_{0}^{2}\right)}{2\left(u_{0}^{2}-c_{0}^{2}\right)^{3/2}}\eta,\end{split}$ (47)
@@ V.2 LINEAR RADIATION
which decreases as a function of the velocity $u_{0}$. This “shortening” of the wavelength can be seen when we compare the wave patterns for $M=2$ in Fig. 3 with the other one for $M=5$ in Fig. 4. If we increase $M$ even more, then we observe in numerical simulations a “continuous” density distribution in the region of for small $\eta$. This means that the wavelength becomes here smaller than the size of the obstacle, so the interference of generated harmonics becomes destructive and it does not lead to formation of a regular wake wave pattern.
@@
Figure 4: Numerical simulation performed for a uniform quasi-2D system with an impenetrable obstacle of radius $R=1$, evolution time $t=25$, and Mach number $M=5$. The density profile is shown in (a), and the corresponding phase in (b).
V.3 OBLIQUE DARK SOLITON AND ITS STABILITY
@@
The oblique dark solitons produced by the LHY fluid inside the Mach cone attains a stationary form only when the critical flow velocity is reached. Below this critical velocity, the soliton is unstable and, similar to the case of the Gross-Pitaevskii equation, it breaks into vortex-antivortex pairs due to the snake instability. A discussion of this instability in the context of the Gross-Pitaevskii equation can be found in Ref. [26] and qualitatively similar behavior is expected for the case of LHY dynamics. An example of a non-stationary street of vortices instead of a stationary oblique soliton is shown in Fig. 3. Above the critical velocity $M>M_{c}$, the oblique dark soliton is formed, as is shown in Fig. 4. A rough estimate based on our numerical results suggests that the critical Mach number for soliton formation lies in the range $3\lesssim M_{c}\lesssim 3.5$. This value is more than twice the critical Mach number for oblique soliton formation in the dimensionless Gross–Pitaevskii equation, $M_{c}^{\text{GPE}}\approx 1.46$, as calculated in Ref. [26]. Below this critical flow velocity, the oblique solitons are absolutely unstable and even for supersonic velocities $1<M<M_{c}$ vortex streets are generated instead. For subsonic flow velocities above $M_{1}\approx\sqrt{2/11}\approx 0.43$, as was shown in Ref. [27], generation of vortex streets leads to loss of superfluidity. Usually, these vortex streets consist of vortex pairs with opposite polarities. However, as was shown in Ref. [28], in case of small values of the obstacle radius $0.04<R<0.13$ there is a region of velocities $0.4\lesssim M\lesssim 0.56$ where Bénard-von Kármán-like streets of vortex pairs with the same polarity are generated. Obviously, these regions of parameters are beyond the region of generation of oblique solitons which we are interested in.
There is very good agreement between the analytical solution for the soliton density profile, given by Eq. (45) for stable oblique solitons, and the numerical profile extracted from Fig. 4 at x=80x=-80, as demonstrated in Fig. 5.
[…]
$\alpha(\rho_{m})=\arccos\left(\frac{1}{|u_{0}|}\sqrt{\frac{2\rho_{m}\left(\frac{3}{5}\rho_{0}^{5/2}+\frac{2}{5}\rho_{m}^{5/2}-\rho_{0}^{3/2}\rho_{m}\right)}{(\rho_{m}-\rho_{0})^{2}}}\right).$ (50)
This expression shows that although both the soliton depth $\rho_{m}$ and the oblique soliton angle $\alpha$ implicitly depend on the obstacle size, this dependence does not affect the accuracy of the theoretical prediction. This can be observed in Fig. 6, where the pair $(\rho_{m},\alpha)$ was obtained varying the value of the obstacle size in the range $0.1\leq R\leq 1.5$. However, for sufficiently large obstacles ($R\gg 1$), the number of oblique soliton pairs formed behind the obstacle increases, and the theory should then be applied to each soliton individually.
@@
Figure 5: Analytical and numerical profiles of the dark soliton for $M=5$ and x=80x=-80. The analytical results were obtained using Eq. (45), where the value of the angle $\alpha=85.035^{\circ}$ was found from the numerical calculations.
Figure 6: Analytical and numerical results for the minimum soliton density $\rho_{m}$ and the oblique soliton angle $\alpha$. The analytical results are given by Eq. (50). The numerical points are obtained by varying the obstacle radius $0.1\leq R\leq 1.5$.
@@ VI Conclusion
As proposed in Ref. [29], the framework of a flow past a barrier serves as an experimental tool to measure and analyze critical velocities associated with the emission of collective excitations. The theoretical approach developed here should be useful to understand the behaviour of linear waves and solitons generated by the interaction of strong laser beams—approximated by an impenetrable barrier—with quasi-2D supersonic quantum droplets or strongly interacting mixtures, where LHY corrections dominate over the mean-field term.
As a direction for future work, we propose extending this approach to the case of supersonic flow past weak, penetrable obstacles, where the local flow velocity varies along the soliton line. As demonstrated in Ref. [30] for a cigar-shaped BEC, the properties of the obstacle also influence the behavior and stability of resulting excitations.